# PRISM: Robust VLM Alignment with Principled Reasoning for Integrated Safety in Multimodality

### v1

@@ Abstract
Safeguarding vision-language models (VLMs) is a critical challenge, as existing methods often suffer from over-defense, which harms utility, or rely on shallow alignment, failing to detect complex threats that require deep reasoning. To this end, we introduce PRISM (Principled Reasoning for Integrated Safety in Multimodality), a system2-like framework that aligns VLMs by embedding a structured, safety-aware reasoning process. Our framework consists of two key components: PRISM-CoT, a dataset that teaches safety-aware chain-of-thought reasoning, and PRISM-DPO, generated via Monte Carlo Tree Search (MCTS) to further refine this reasoning through Direct Preference Optimization to help obtain a delicate safety boundary. Comprehensive evaluations demonstrate PRISM’s effectiveness, achieving remarkably low attack success rates including 0.15% on JailbreakV-28K for Qwen2-VL and 90% improvement over the previous best method on VLBreak for LLaVA-1.5. PRISM also exhibits strong robustness against adaptive attacks, significantly increasing computational costs for adversaries, and generalizes effectively to out-of-distribution challenges, reducing attack success rates to just 8.70% on the challenging multi-image MIS benchmark. Remarkably, this robust defense is achieved while preserving, and in some cases enhancing, model utility. To promote reproducibility, we have made our code, data, and model weights available at https://github.com/SaFoLab-WISC/PRISM.
@@ 1 Introduction
Figure 1: Helpfulness and harmlessness trade-off on LLaVA-1.5 and Qwen2-VL.
@@
The rapid advancement of Vision-Language Models (VLMs) Liu et al. (2023); Wang et al. (2024b); Zhu et al. (2025) has shown great potential in tasks ranging from image understanding Wang et al. (2024a) and visual question answering Yu et al. (2024); Liu et al. (2024) to complex multimodal reasoning Lu et al. (2024), positioning them as foundational technologies for next-generation AI applications. However, the multimodal nature of these models introduces unprecedented challenges in safety alignment, as malicious actors can exploit vulnerabilities through carefully crafted text, image, or combinations of text and images to lead the model to bypass the internal safety alignment and generate unsafe answers Luo et al. (2024); Wang et al. (2024e; c).
@@
As illustrated in Figure 3, we identify three distinct categories of safety violations in multimodal contexts: (1) Problem unsafe, where the textual prompt contains explicitly harmful content; (2) Image unsafe, where the visual input presents safety risks or violates content policies; and (3) Problem+Image combination unsafe, where neither the text nor image is inherently harmful, but their combination creates safety concerns. This third category represents a particularly challenging threat model, as it exploits the complex semantic relationships between modalities to embed malicious intent within seemingly innocent multimodal inputs, making them invisible to single-modality alignment strategies.
@@
Figure 2: Response comparison between existing defense method with our proposed PRISM method.
@@
Existing defenses typically rely on heuristic safeguards—such as simple keyword filters, visual classifiers, or static refusal rules. These strategies are effective against overtly unsafe text or images, but falter in subtle cross-modal attacks where harm only emerges through the interaction between modalities. The efficacy of training-based defense mechanisms, such as those utilizing Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2025b); Ji et al. (2025) or Supervised Fine-Tuning (SFT) Zong et al. (2024), is often constrained by their tendency to produce rote refusal responses. A significant limitation of these approaches is their lack of detailed reasoning, which fails to explicitly articulate or reveal the malicious intent underlying a prompt. Consequently, as illustrated in Figure 2, such methods often yield shallow alignment Qi et al. (2025), where they fail to develop a genuine understanding of safety principles and instead rely on superficial pattern matching that can be easily circumvented by sophisticated attacks.
To address this, we take inspiration from System 2-style reasoning that analyzes the interplay between textual and visual elements. Inspired by the success of reasoning to protect large language models Zhang et al. (2025a), effective safety mechanisms should employ chain-of-thought reasoning that carefully examines the context of the problem, analyzes visual content relative to queries, and synthesizes multimodal information to identify safety concerns.
To this end, we introduce PRISM, a safety alignment framework that leverages structured reasoning processes to enhance VLM safety while minimally compromising utility. Our approach consists of two key components: PRISM-CoT, a curated dataset that demonstrates proper safety-aware chain-of-thought reasoning through four structured stages where Problem aims to identify potential harmful content in textual prompts, Caption provides contextualized visual understanding, Reasoning synthesizes multimodal information to detect safety violations, and Output generates appropriate responses with explicit safety justification. PRISM-DPO, a preference optimization dataset generated through Monte Carlo Tree Search (MCTS) that contains step-level preference pairs for direct preference optimization Rafailov et al. (2023).
Our comprehensive evaluations demonstrate that PRISM-DPO significantly enhances the safety of Vision-Language Models (VLMs) while preserving their core utility. The model exhibits robust safety across diverse attack vectors. On static benchmarks like JailbreakV-28K and VLBreakBench, it achieves a near-zero Attack Success Rate (ASR). It also displays strong resilience against adaptive attacks, successfully defending against repeated queries on models like LLaVA-1.5 and thus substantially increasing the attacker’s cost. Furthermore, our framework generalizes effectively to out-of-distribution challenges, achieving a low ASR ($\sim 5\%$) on the multi-image MIS benchmark. Critically, these safety gains do not compromise helpfulness, as our model attains a state-of-the-art score on the MM-Vet-v2 benchmark. As illustrated in Figure 1, our approach strikes an optimal balance between safety and utility, showing that structured reasoning and preference optimization can work synergistically. Finally, our results indicate that safety can be further amplified at test-time by scaling the reasoning framework, highlighting its effectiveness and potential.
@@ 2 Related Work
Jailbreaking Vision Language Models. The landscape of Vision-Language Model (VLM) jailbreaking is evolving from simple structural exploits Liu et al. (2024); Luo et al. (2024) to sophisticated attacks that require complex multimodal reasoning. Initial research exposed vulnerabilities through static strategies like Figstep Gong et al. (2025), FC-Attack Zhang et al. (2025c), and MML Wang et al. (2024e), which used crafted input combinations, as well as adaptive frameworks like IDEATOR Wang et al. (2024c). However, a more profound shift is that superficial safety alignments are no longer sufficient, prompting the development of a new generation of benchmarks to probe deeper reasoning vulnerabilities. For instance, MSSBench Zhou et al. (2025) assesses the safety of a language query within its visual context, demanding contextual understanding rather than simple policy adherence. Similarly, VLSBench Hu et al. (2024) requires multimodal cooperation to succeed, focusing on revealing subtle safety cue leakages in text. Furthermore, MIS Ding et al. (2025b) integrates multi-image inputs to test a model’s advanced visual reasoning in complex scenarios. Collectively, these benchmarks signify that the frontier of VLM safety has moved beyond shallow alignment, demanding models with a fundamentally more robust and context-aware reasoning capability.
Safeguarding Vision Language Models. Approaches to safeguarding Vision-Language Models (VLMs) can be broadly categorized into training-based and training-free methods Ding et al. (2025a); Pi et al. (2024). Focusing on training-based approaches, current work primarily follows two paradigms: Supervised Fine-Tuning (SFT) on safety datasets, as seen in Dress Chen et al. (2024b) and VLGuard Zong et al. (2024), and preference optimization using Direct Preference Optimization (DPO), as employed by SPA-VL Zhang et al. (2025b) and SafeRLHF-V Ji et al. (2025). However, by training models for refusal without an explicit reasoning process, these methods often lead to two shortcomings: over-defense, where models incorrectly reject benign queries, and shallow alignment, which leaves them vulnerable to sophisticated attacks that require reasoning to uncover concealed harmful intent.
@@ 3 Safety-Aware Reasoning Dataset Generation
To be precise, we curate a structured reasoning safety dataset with the format of four CoT reasoning steps: Problem, Caption, Reasoning, and Output.
@@ 3.1 Dataset Preparation
Our dataset preparation methodology follows a systematic approach to curating a comprehensive multimodal safety dataset that encompasses the full spectrum of safety violations requiring chain-of-thought reasoning.
Problem-Image Combination Safety Violation Generation. While existing datasets primarily focus on violations at the individual modality level (problem-unsafe or image-unsafe), we recognize a critical gap in addressing combination-unsafe scenarios where neither the text nor image is inherently harmful, but their interaction creates safety concerns requiring sophisticated reasoning. To address this limitation, we adopt the methodology from MML Wang et al. (2024e), which employs steganographic techniques including word replacement, base64 encoding, and rotation to embed malicious content within seemingly benign image-text pairs. This approach creates about 3k instances where the harmful intent is only revealed through careful analysis of the multimodal interaction, necessitating advanced reasoning capabilities for detection.
@@ 3.2 Safety-aware Chain-of-Thought Reasoning Process Generation
Drawing upon insights from AdaShield Wang et al. (2024d), we recognize the importance of explicitly specifying response methodologies in defense prompts by clearly identifying the specific type of violation that occurs. To this end, we leverage the safety categories annotated in the original datasets and systematically map them to eight distinct safety violation types as defined by Wang et al. (2025). With the given safety categories, we generate a structured chain-of-thought reasoning process that includes four key steps: Problem, Caption, Reasoning, and Output with refusal response once detecting malicious intent, and outputs a refusal that points out the specific type of safety violation.
We employ GPT-4o to generate structured chain-of-thought reasoning processes through a systematic four-step framework:
PROBLEM. This initial step entails a comprehensive examination of the textual prompt to identify potential harmful content or malicious intent. We provide only the text prompt and instruct it to analyze the underlying intention behind the query and determine whether any safety violations are present. Upon detection of a violation, GPT-4o explicitly identifies the specific category of safety infringement according to our established taxonomy.
CAPTION. Recognizing that effective image safety evaluation requires contextual understanding, we instruct GPT-4o to generate a caption that describes the visual content specifically in relation to the problem context. This contextualized captioning approach facilitates the identification of safety violations that might be apparent only when considering the image within the framework of the accompanying query.
REASONING. This critical step involves the integration and synthesis of information derived from both the problem analysis and visual assessment. GPT-4o conducts a systematic evaluation of the interaction between textual and visual modalities, identifying potential safety concerns that emerge specifically from their combination, which is a capability essential for detecting sophisticated multimodal safety violations.
OUTPUT. Upon detection of any safety violation, GPT-4o generates an appropriate refusal response that explicitly articulates the specific type of violation identified in the preceding analytical steps. This ensures alignment with established safety protocols while maintaining transparency regarding the reasoning behind the refusal.
This structured reasoning framework ensures coverage of potential safety violations across individual modalities and their interactions. When malicious intent is detected, subsequent reasoning steps maintain the structural integrity of the process while appropriately refusing to engage with harmful content, thereby creating valuable training examples for our fine-tuning procedure. For a detailed exposition of the prompting methodology, please refer to Appendix B.
@@
Following the generation process, we implement a quality check to verify proper formatting and filter out any jailbroken responses. This yields just under 3,000 high-quality safety reasoning instances. Crucially, this generation process was not unconstrained; we leveraged the carefully crafted safety violation labels from the original datasets as explicit guidance for GPT-4o. By providing these ground-truth labels, we anchor the generated reasoning to human-annotated safety standards, thus ensuring the process is not solely limited by the intrinsic capabilities of the generator model. To preserve performance on benign tasks and prevent over-refusal, we then augment this safety-focused data with 3,000 benign instances from LLaVA-CoT, adapted to our four-step format. This balanced composition results in a principled supervised fine-tuning dataset of approximately 6k instances, which we designate as PRISM-CoT.
@@
Figure 4: Overview of our safety-aware MCTS preference data generation process. (a) Illustrates an image-unsafe instance example, where safety rewards are computed by GPT-4o without back-propagation. (b) Demonstrates a benign instance example, where helpfulness rewards are calculated by GPT-4o using existing reasoning steps as evaluation criteria, with rewards back-propagated through the decision tree.
@@ 3.3 Safety-aware MCTS Preference Generation
where $\epsilon$ is a small margin to ensure statistical significance and $\theta$ is the threshold for an accepted node.
Following the MCTS preference generation methodology described above, we initiate the process with a carefully curated dataset comprising 500 benign and 1,500 malicious image-text pairs to systematically generate reasoning steps for safety-aware reasoning. Our implementation employs $k=3$ candidate children for each node during the expansion phase, with a maximum iteration limit of $200$ per reasoning tree to ensure computational tractability while maintaining sufficient exploration depth.
For preference pair generation, we establish a difference margin of $\epsilon=0.4$ and a quality threshold of $\theta=0.8$ to identify statistically significant reward differences between reasoning paths. Recognizing that each non-terminal node contributes meaningful insights to the overall reasoning process, we extract preference pairs across all reasoning levels rather than limiting collection to terminal nodes.
Through this systematic preference generation procedure, we obtain a total of 10,000 preference pairs, which we designate as PRISM-DPO. This substantial preference dataset serves as the foundation for our subsequent direct preference optimization training phase.
@@ 4 Experiment
4.1 Supervised Fine-tuning on PRISM-CoT
As illustrated in Figure 3, we fine-tune our models on the PRISM-CoT dataset, with each reasoning step delimited by specialized tokens. We augment the model vocabulary with eight special tokens to demarcate the structured reasoning process. The fine-tuning procedure involves comprehensive parameter optimization using $90\%$ of the dataset for training and $10\%$ for validation. We conduct training with a batch size of $16$, a learning rate of 1e51e-5, and train for $5$ epochs. All experiments are performed on $4$ NVIDIA A100 (40GB) GPUs, with each GPU processing $4$ samples per batch.
4.2 Direct Preference Optimization on PRISM-DPO
Following the supervised fine-tuning phase, we proceed with direct preference optimization (DPO) training on the PRISM-DPO dataset. This phase involves optimizing the model to achieve finer-grained alignment at the reasoning step level. For DPO implementation, we employ LoRA fine-tuning with rank $r=16$ and scaling factor $\alpha=64$. To stabilize the training process, we incorporate a mixture of DPO and SFT objectives with a ratio of $1.0$. We utilize a batch size of $16$ and a learning rate of 1e51e-5, conducting training for $3$ epochs across $4$ NVIDIA A100 (40GB) GPUs, with each GPU processing $4$ samples per batch.

### v2

@@ Abstract
Safeguarding vision-language models (VLMs) is a critical challenge, as existing methods often suffer from over-defense, which harms utility, or rely on shallow alignment, failing to detect complex threats that require deep reasoning. To this end, we introduce PRISM (Principled Reasoning for Integrated Safety in Multimodality), a System 2-like framework that aligns VLMs through a structured four-stage reasoning process explicitly designed to handle three distinct categories of multimodal safety violations. Our framework consists of two key components: a structured reasoning pipeline that analyzes each violation category in dedicated stages, and PRISM-DPO, generated via Monte Carlo Tree Search (MCTS) to refine reasoning quality through Direct Preference Optimization. Comprehensive evaluations show that PRISM substantially reduces attack success rates on JailbreakV-28K and VLBreak, improves robustness against adaptive attacks, and generalizes to out-of-distribution multi-image threats, while better preserving model utility on benign multimodal benchmarks. Our code, data, and model weights available at https://github.com/SaFoLab-WISC/PRISM.
PRISM: Robust VLM Alignment with Principled Reasoning
for Integrated Safety in Multimodality
Nanxi Li1, Zhengyue Zhao1, G. Edward Suh2,4, Marco Pavone3,4, Chaowei Xiao1,4 1Johns Hopkins University, 2Cornell University, 3Stanford University, 4NVIDIA
@@ 1 Introduction
The rapid progress of Vision-Language Models (VLMs) Liu et al. (2023); Wang et al. (2024b); Zhu et al. (2025) has unlocked strong capabilities across a wide range of tasks, including image understanding Wang et al. (2024a), visual question answering Yu et al. (2024); Liu et al. (2024), and complex multimodal reasoning Lu et al. (2024). These advances position VLMs as foundational technologies for next-generation AI applications. However, their multimodal nature also introduces new and pressing challenges for safety alignment.
Figure 1: Performance of different methods by using LLaVA-1.5 and Qwen2-VL as base models. Our method achieves a better Helpfulness and harmlessness trade-off
@@
While perturbation-based adversarial attacks have been extensively studied Wu et al. (2021); Qi et al. (2024) and can often be mitigated through purification methods Guo et al. (2024) or adversarial training Qi et al. (2024); Cui et al. (2024), structure-based jailbreak attacks Luo et al. (2024); Wang et al. (2025a, 2024c) remain an open challenge. In such attacks, adversaries craft harmful text, images, or text–image combinations to bypass alignment mechanisms and elicit unsafe outputs Luo et al. (2024); Wang et al. (2025a, 2024c). As illustrated in Figure 3, there are three distinct categories of safety violations in multimodal contexts: (1) Problem unsafe, where the textual prompt contains explicitly harmful content; (2) Image unsafe, where the visual input presents safety risks or violates content policies; and (3) Problem+Image combination unsafe, where neither the text nor image is inherently harmful, but their combination creates safety concerns. This third category represents a challenging threat model, as it exploits the complex semantic relationships between modalities to embed malicious intent within seemingly innocent multimodal inputs, making them invisible to single-modality alignment strategies.
@@
Figure 2: Response comparison between the existing defense method with our proposed PRISM method.
@@
Existing defenses typically rely on heuristic safeguards—such as simple keyword filters, visual classifiers, or static refusal rules. These strategies are effective against overtly unsafe text or images, but falter in subtle cross-modal attacks where harm only emerges through the interaction between modalities. The efficacy of training-based defense mechanisms, such as those utilizing Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2025b); Ji et al. (2025) or Supervised Fine-Tuning (SFT) Zong et al. (2024), is often constrained by their tendency to produce rote refusal responses. A significant limitation of these approaches is their lack of detailed reasoning, which fails to explicitly reveal the malicious intent underlying a prompt. Consequently, as illustrated in Figure 2, such methods often yield shallow alignment Qi et al. (2025), where they fail to develop a genuine understanding of safety principles and instead rely on superficial pattern matching that can be easily circumvented by sophisticated attacks.
To address this, we introduce PRISM, a safety alignment framework that moves beyond shallow refusal. The non-triviality of multimodal safety reasoning lies in the fact that neither the text nor the image alone may trigger a safety violation; it is only through their semantic interaction that malicious intent is revealed. PRISM addresses this via two key components:
PRISM-CoT: A curated dataset demonstrating four-stage reasoning (Problem, Caption, Reasoning, Output). This structure is specifically designed to isolate visual context from textual intent, a critical step for detecting combination-unsafe attacks that exploit cross-modal semantic gaps.
PRISM-DPO: A step-level preference optimization dataset generated via Monte Carlo Tree Search (MCTS). This is a principled approach to credit assignment in safety: we initialize the search from the SFT model and apply a stage-local reward strategy. This ensures that a model which reasons correctly about a context is not unfairly penalized for a later refusal or decision error, which is a key driver behind our improved safety–utility trade-off.
Our comprehensive evaluations show that PRISM-DPO substantially improves robustness against multimodal jailbreak attacks while retaining strong performance on benign multimodal tasks. On static benchmarks such as JailbreakV-28K and VLBreakBench, PRISM significantly reduces Attack Success Rate (ASR) compared to the undefended base models and prior training-based defenses. It also improves resilience under adaptive attacks, increasing the number of attacker queries required to find a successful jailbreak. Moreover, PRISM generalizes to out-of-distribution multi-image threats on MIS with consistently lower ASR than competing defenses. Importantly, these safety gains come with markedly better utility retention than existing alignment baselines: across base models, PRISM still maintains good MM-Vet-v2 helpfulness, indicating an improved safety–utility trade-off (Figure 1). Finally, we show safety can be further strengthened via test-time scaling of the structured reasoning framework.
@@ 2 Related Work
Jailbreaking Vision Language Models. The landscape of Vision-Language Model (VLM) jailbreaking is evolving from simple structural exploits Luo et al. (2024) to sophisticated attacks that require complex multimodal reasoning. Initial research exposed vulnerabilities through static strategies like Figstep Gong et al. (2025), FC-Attack Zhang et al. (2025c), and MML Wang et al. (2025a), which used crafted input combinations, as well as adaptive frameworks like IDEATOR Wang et al. (2024c). However, a more profound shift is that superficial safety alignments are no longer sufficient, prompting the development of a new generation of benchmarks to probe deeper reasoning vulnerabilities. For instance, MSSBench Zhou et al. (2024) assesses the safety of a language query within its visual context, demanding contextual understanding rather than simple policy adherence. Similarly, VLSBench Hu et al. (2025) requires multimodal cooperation to succeed, focusing on revealing subtle safety cue leakages in text. Furthermore, MIS Ding et al. (2025) integrates multi-image inputs to test a model’s advanced visual reasoning in complex scenarios. Collectively, these benchmarks signify that the frontier of VLM safety has moved beyond shallow alignment, demanding models with a fundamentally more robust and context-aware reasoning capability.
Safeguarding Vision Language Models. Approaches to safeguarding Vision-Language Models (VLMs) can be broadly categorized into training-based and training-free methods Ding et al. (2024); Pi et al. (2024). Focusing on training-based approaches, current work primarily follows two paradigms: Supervised Fine-Tuning on safety datasets, as seen in Dress Chen et al. (2024b) and VLGuard Zong et al. (2024), and preference optimization using Direct Preference Optimization, as employed by SPA-VL Zhang et al. (2025b) and SafeRLHF-V Ji et al. (2025). However, by training models for refusal without an explicit reasoning process, these methods lead to two shortcomings: over-defense, where models incorrectly reject benign queries, and shallow alignment, which leaves them vulnerable to sophisticated attacks that require reasoning to uncover concealed intent.
Reasoning-based Safety Alignment. Recent work demonstrates that embedding explicit reasoning into the alignment process yields more robust safety than shallow refusal training. In the text-only domain, STAIR Zhang et al. (2025a) applies introspective reasoning with MCTS-based preference optimization to improve LLM safety alignment. Deliberative alignment Guan et al. (2024) shows that training models to explicitly reason over safety specifications before responding substantially improves jailbreak robustness while reducing over-refusal. Process-level supervision Lightman et al. (2023) has proven more effective than outcome-only supervision for verifying reasoning chains. Self-rewarding language models Yuan et al. (2024) demonstrate that models can serve as their own reward signals for iterative improvement. While these works establish the value of reasoning for text-only safety, extending this paradigm to the multimodal domain introduces fundamentally distinct challenges: VLMs must jointly reason over visual and textual inputs to detect cross-modal threats where neither modality alone reveals harmful intent. PRISM addresses this gap by designing a multimodal-specific structured reasoning framework with stage-local reward assignment tailored to the unique threat model of vision-language interactions.
@@ 3.1 Dataset Preparation
Our dataset preparation methodology follows a systematic approach to curating a comprehensive multimodal safety dataset that encompasses the full spectrum of safety violations requiring chain-of-thought reasoning. For details about the dataset selection, please refer to the Appendix B.
Problem-Image Combination Safety Violation Generation. While existing datasets primarily focus on violations at the individual modality level (problem-unsafe or image-unsafe), we recognize a critical gap in addressing combination-unsafe scenarios where neither the text nor image is inherently harmful, but their interaction creates safety concerns requiring sophisticated reasoning. To address this limitation, we adopt the methodology from MML Wang et al. (2025a), which employs steganographic techniques including word replacement, base64 encoding, and rotation to embed malicious content within seemingly benign image-text pairs. This approach creates about 3k instances where the harmful intent is only revealed through careful analysis of the multimodal interaction, necessitating advanced reasoning capabilities for detection.
@@ 3.2 Safety-aware Chain-of-Thought Reasoning Process Generation
Drawing upon insights from AdaShield Wang et al. (2024d), we recognize the importance of explicitly specifying response methodologies in defense prompts by clearly identifying the specific type of violation. We leverage the safety categories annotated in the original datasets and systematically map them to eight distinct safety violation types as defined by Wang et al. (2025b). We then generate structured chain-of-thought reasoning processes through a four-step framework: PROBLEM (analyzing textual intent and identifying safety violations), CAPTION (generating contextualized visual descriptions relative to the query), REASONING (synthesizing cross-modal information to detect multimodal safety concerns), and OUTPUT (producing appropriate refusals with explicit violation justification or helpful answers for benign queries). Detailed descriptions of each step and the prompts are provided in Appendix C.
@@
Crucially, this generation process is grounded in the original datasets rather than relying on the curator model to infer safety intent: for each instance, we provide the human-annotated safety category (and its definition/policy specification) from the source benchmark as explicit conditioning, and require the generated rationale and final decision to be consistent with that label. Following generation, we apply formatting and consistency checks and filter any traces that exhibit unsafe compliance or label-inconsistent decisions, yielding 3,000 high-quality safety reasoning instances. To preserve performance on benign tasks and prevent over-refusal, we augment this safety-focused data with 3,000 benign instances from LLaVA-CoT adapted to our four-step format. This balanced composition results in a principled supervised fine-tuning dataset of approximately 6k instances, which we designate as PRISM-CoT.
@@
Figure 4: Overview of our safety-aware MCTS preference data generation process. (a) Illustrates an image-unsafe instance example, where safety rewards are computed by a judger without back-propagation. (b) Demonstrates a benign instance example, where helpfulness rewards are assigned by a judger using existing reasoning steps as evaluation criteria, with rewards back-propagated through the decision tree.
@@ 3.3 Safety-aware MCTS Preference Generation
where $\epsilon$ is a margin to ensure statistical significance and $\theta$ is the threshold for an accepted node.
We extract preference pairs across all reasoning levels rather than limiting collection to terminal nodes. Through this procedure, we obtain a total of 10,000 preference pairs, which we designate as PRISM-DPO. Detailed hyperparameters for MCTS generation are provided in Appendix F.